# Training-free Anomaly Event Detection via LLM-guided Symbolic Pattern Discovery

### v2

@@ 3.2 Automated LLM Reasoning
To enhance the efficiency and effectiveness of symbolic pattern discovery, we propose an automated reasoning mechanism that leverages the semantic understanding capabilities of Large Language Models (LLMs). This section details our structured prompt space design and its integration with symbolic logic search.
@@ 3.2.1 Structured Prompt Space
$P_{init}=\{\text{scene},\text{entities},\text{constraints}\}$ (7)
This activation prompt triggers the LLM’s prior knowledge relevant to the specific anomaly detection scenario, creating crucial connections between visual entities and semantic understanding.
@@
$P_{cot}=\{s_{1}\rightarrow s_{2}\rightarrow...\rightarrow s_{k}\}$ (8)
where each step $s_{i}$ guides the LLM through professional analytical frameworks for identifying potential anomaly patterns. This systematic approach ensures comprehensive consideration of entity relationships and domain constraints.
@@
$P_{feed}=\{(r_{1},\alpha_{1}),(r_{2},\alpha_{2}),...,(r_{n},\alpha_{n})\}$ (9)
@@ 3.2.2 LLM-Guided Symbolic Search
This bidirectional interaction accelerates the discovery of meaningful anomaly patterns while maintaining the interpretability of the detection process. The LLM’s suggestions help navigate the vast space of possible symbolic expressions, while the symbolic search framework ensures that the final patterns remain explicit and verifiable.
Analysis. Through the integration of symbolic logic search and LLM reasoning, our approach offers several distinct advantages. First, the training-free nature of our framework eliminates the need for extensive labeled datasets, as the symbolic pattern discovery process operates directly on entity-level features extracted from test samples. This characteristic is particularly valuable in real-world scenarios where collecting and annotating training data is costly or impractical. Second, our method achieves high interpretability through two complementary mechanisms: (a) the discovered symbolic expressions provide explicit, human-readable logical patterns that directly explain the detection decisions, and (b) the LLM’s reasoning process offers semantic context for these patterns, making them more accessible to domain experts. This dual-layer interpretability sets our approach apart from traditional black-box models. Third, the bidirectional interaction between symbolic search and LLM reasoning creates an efficient optimization process. The LLM’s semantic understanding and common sense knowledge help navigate the vast space of possible symbolic expressions, significantly reducing the search complexity compared to pure evolutionary approaches. Meanwhile, the symbolic search framework provides objective evaluation criteria that prevent the accumulation of LLM reasoning errors, ensuring the robustness of the final patterns. Lastly, our structured prompt space design enables systematic utilization of LLM capabilities while maintaining consistency in the reasoning process. This structured approach not only improves the quality of LLM suggestions but also makes the entire framework more stable and reproducible compared to methods relying on ad-hoc prompting strategies.
@@ 4 Experiment
Table 1: Performance of different open set detectors on multiple data sets with or without VED-SR module. (AUROC%)
@@
Original +VED-SR Original +VED-SR Original +VED-SR
@@ 4.1 Experimental Setup
To comprehensively evaluate our approach, we conduct experiments on four diverse datasets: two large-scale self-collected datasets and two public benchmarks, each targeting different aspects of anomaly event detection.
@@ 4.1.1 Self-collected Datasets
Multi-Event Dataset is our large-scale dataset about detecting multiple anomaly event in complex scenarios. The dataset comprises over 110,000 images spanning various scenarios including fires and waste incineration (30,954 images), multi-rods fishing (12,000 images), night fishing (97 images), license plate detection (12,487 images), personnel loitering and intrusion (10,788 images), among others. For this study, we specifically focus on the multiple-rod fishing scenario, where we have 15,000 training images with 45,341 detailed bounding box annotations covering persons, fishing rods, tackle bags, and umbrellas. The test set contains 2,283 images, with 1,098 cases showing anomalous multi-rods fishing activities and 1,185 normal cases with single-rod fishing. The rich annotations enable precise analysis of entity relationships and logical constraints.
@@ 4.1.2 Public Benchmarks
ERA Dataset [35] (Event Recognition in Aerial videos) provides a comprehensive collection of aerial footage covering various event categories. We organize our evaluation around three main event categories: BALL events (327 images) encompassing baseball, soccer, and basketball games; Person_crowded events (352 images) including concerts, conflicts, parade protests, and parties; and Sport events (258 images) covering cycling, boating, and racing activities. Additionally, we utilize 347 Non-event images as negative samples, creating a balanced evaluation framework for our method’s discriminative capabilities across different event types.
@@ 4.1.3 Implementation Details
Our implementation consists of three main components: multimodal detection, symbolic regression, and computational infrastructure. We detail each aspect to ensure reproducibility.
Multimodal Detection Setup We employ three state-of-the-art multimodal detectors in our experiments: APE [46] (serving as the primary detector), GLIP [26], and YOLO-WORLD [10]. To optimize detection performance, we implement a two-stage prompt generation process. Initially, we leverage LLM to analyze event scenario descriptions and generate comprehensive detection prompts. The LLM not only generates prompts for objects directly related to the event scenarios but also produces prompts for contextually related but non-event objects, ensuring broad coverage of potential scene elements. Based on empirical studies of detector characteristics, we configure different detection thresholds for optimal performance. Considering APE’s characteristically lower threshold nature, we set its minimum detection threshold to 0.05. For GLIP and YOLO-WORLD, we establish a higher threshold of 0.1 to maintain a balance between precision and recall in object detection.
@@
Computational Resources Our experimental framework is deployed on a cluster of 8 NVIDIA V100 GPUs, each equipped with 32GB memory. This computational infrastructure supports three main tasks: the inference operations of base object detectors, the deployment and prompt processing of Large Language Models, and the computational requirements of symbolic regression optimization. This hardware configuration ensures efficient processing of our training-free anomaly detection pipeline while maintaining real-time performance capabilities.
Table 2: The overall performance on the UCSD ped2 [52] benchmark.This intuitively reflects that the performance of VED-SR (APE) without training is very close to the current SOTA.
@@
VED-SR 98.7
4.2 Main Results
We conduct comprehensive experiments to evaluate our approach from multiple perspectives: effectiveness of the training-free framework, necessity of fine-tuning, and comparison with traditional methods.
Table 3: Performance of different fine-tuned methods and with or without VED-SR module. Lora indicates whether to Lora-tuning the APE model. Prompt indicates whether to Prompt-tuning the APE model (AUROC%)
VED-SR Lora Prompt Helmet-Mac Multi-rods-Fishing
67.41 66.82
83.18 (+15.77) 75.16 (+8.34)
84.11 75.86
95.67 (+11.56) 78.44 (+2.58)
67.42 66.82
81.62 (+14.2) 76.06 (+9.24)
Comparison with Different Detection Architectures. We first evaluate our framework using three state-of-the-art open-vocabulary detectors: APE, GLIP, and YOLO-WORLD, which represent diverse architectural choices in both visual and language processing. These detectors employ different visual backbones (VIT [14], Swin-L [34], and YOLOv8) and language models (EVA-CLIP [50], CLIP [39], and BERT [12]). As shown in Table 1, all three detectors achieve strong performance without any fine-tuning, with APE consistently outperforming others across all five anomaly event scenarios. This superior performance of APE can be attributed to its more sophisticated visual-language alignment mechanism and larger pre-training dataset, which enables better transfer of knowledge to anomaly detection tasks. The consistent performance across architecturally diverse models also suggests that our framework’s effectiveness is not tied to specific architectural choices, but rather stems from the fundamental synergy between symbolic reasoning and detection capabilities. Finding 1: The effectiveness of our training-free framework is architecture-agnostic, with APE’s superior performance likely due to its enhanced visual-language alignment and broader pre-training.
Comparison with Fine-tuning Approaches. To further validate the efficiency of our training-free approach, we conduct comparative experiments with fine-tuned variants on the Helmet-Mac and Multi-rods Fishing datasets. We implement two common fine-tuning strategies: LORA fine-tuning[17] and Prompt tuning [25], representing different levels of parameter adaptation. Results in Table 3 demonstrate that our training-free approach achieves comparable performance to fine-tuned models. This intriguing finding suggests that our training-free approach effectively leverages the model’s general understanding of visual-language relationships. Additionally, the symbolic reasoning component provides a more structured way to capture anomaly patterns compared to implicit learning through fine-tuning, achieving similar effectiveness without the computational overhead of parameter adaptation. Finding 2: Our training-free approach is comparable to fine-tuning methods, possibly due to better preservation of general visual-language understanding and more structured pattern discovery.
Comparison with Traditional Methods. For comparison with traditional anomaly detection methods, we evaluate our approach on UCSD Ped2 [52] dataset. Our method achieves an impressive 98.7% accuracy without utilizing semantic-level annotations or task-specific fine-tuning, approaching the state-of-the-art performance (99.7%), showed in Table 2. This minimal performance gap is particularly interesting considering the vast difference in approach complexity. We hypothesize that this effectiveness stems from two factors: first, the pre-trained detectors already possess rich semantic understanding that generalizes well to anomaly detection; second, our symbolic reasoning framework effectively translates this semantic knowledge into explicit detection rules, potentially capturing patterns that are similar to those learned by supervised methods but in a more interpretable manner. Finding 3: The near-SOTA performance on UCSD Ped2 suggests that combining pre-trained knowledge with symbolic reasoning can effectively match supervised learning capabilities.
Figure 4: Performance on VED-SR with or without LLM.
4.3 Ablation Study
@@
Impact of LLM Integration. Figure 4 shows further integration of LLM reasoning not only enhances detection accuracy but accelerates the convergence of symbolic regression, showing a synergistic effect between semantic understanding and pattern discovery. A detailed analysis of the convergence process reveals that LLM guidance helps avoid many unnecessary exploration paths in the symbolic search space, effectively pruning unpromising directions based on semantic understanding. Moreover, we observe the quality of discovered patterns improves with LLM guidance, as the patterns tend to be more semantically meaningful and better aligned with human intuition about anomalies. Finding 4: The significant performance gains from each component suggest that effective anomaly detection requires both systematic pattern discovery (through symbolic regression) and semantic guidance (through LLM reasoning).
These ablation results highlight an important insight: while traditional approaches often treat anomaly detection as either a purely statistical or purely semantic task, our results suggest that combining both aspects leads to superior performance. The symbolic regression provides the systematic exploration capability, while LLM reasoning contributes semantic understanding, creating a framework that leverages the strengths of both approaches. Our experiments demonstrate that the combination of training-free operation, symbolic reasoning, and LLM guidance creates a robust and effective framework for anomaly detection, often matching or exceeding the performance of more complex, training-dependent approaches.
5 Conclusion
In this paper, we present VED-SR, a novel approach that addresses visual event detection through symbolic logical reasoning. Our framework makes three key contributions to the field: First, we establish a new paradigm by uniquely combining symbolic regression for logical pattern search with Large Language Models (LLMs) for symbolic reasoning. This integration successfully transforms complex anomaly event patterns into human-readable symbolic expressions while achieving performance comparable to current state-of-the-art methods. Second, we demonstrate the effectiveness of a training-free framework that eliminates the need for task-specific fine-tuning while maintaining competitive performance. This achievement suggests a promising direction for developing more efficient and adaptable event detection systems. Third, we introduce a new benchmark dataset containing over 5,000 annotated samples for visual event detection, providing a valuable resource for future research in this domain. Looking forward, several promising directions emerge for extending this work, including expanding the symbolic logic search space, automating prompt space construction, and integrating with object detection and active learning frameworks. These developments could further advance the field of visual event detection while maintaining the interpretability and efficiency advantages demonstrated in our work.

### v3

@@ 3.2 Automated LLM Reasoning
To enhance the efficiency and effectiveness of symbolic pattern discovery, we propose an automated reasoning mechanism that leverages the semantic understanding capabilities of Large Language Models (LLMs). This LLM-guided approach consists of two main components: a structured prompt space for eliciting effective reasoning and an integrated symbolic search mechanism that combines LLM suggestions with systematic exploration.
@@ 3.2.1 Structured Prompt Space
$P_{\text{init}}=\{\text{scene},\text{entities},\text{constraints}\}$ (8)
This activation prompt triggers the LLM’s prior knowledge relevant to the specific visual event understanding scenario, creating crucial connections between visual entities and semantic understanding.
@@
$P_{\text{cot}}=\{s_{1}\rightarrow s_{2}\rightarrow...\rightarrow s_{k}\}$ (9)
where each step $s_{i}$ guides the LLM through professional analytical frameworks for identifying potential visual event symbolic patterns. This systematic approach ensures comprehensive consideration of entity relationships and domain constraints.
@@
$P_{\text{feed}}=\{(r_{1},\alpha_{1}),(r_{2},\alpha_{2}),...,(r_{n},\alpha_{n})\}$ (10)
@@ 3.2.2 LLM-Guided Symbolic Search
This bidirectional interaction accelerates the discovery of meaningful symbolic patterns while maintaining the interpretability of the detection process. The LLM’s suggestions help navigate the vast space of possible symbolic expressions, while the symbolic search framework ensures that the final patterns remain explicit and verifiable.
Analysis. Through the integration of symbolic logic search and automated LLM reasoning, our framework offers several key advantages for event understanding in object detection: First, our approach produces inherently interpretable results through complementary mechanisms. The discovered symbolic expressions provide explicit, human-readable logical patterns that directly explain recognition decisions, while the LLM’s reasoning process offers semantic context for these patterns. This dual-layer interpretability is critical for applications where understanding the reasoning process is as important as the final decision. Second, the bidirectional interaction between symbolic search and LLM reasoning creates an efficient optimization process. The LLM’s semantic understanding helps navigate the vast space of possible symbolic expressions, while the symbolic search framework grounds the LLM’s suggestions in empirical performance. Our structured prompt design ensures systematic utilization of LLM capabilities while maintaining consistency and reproducibility in the reasoning process. Finally, our framework provides significant practical deployment advantages. By operating on the outputs of existing object detectors, it eliminates the need for large-scale training datasets typically required by deep learning methods. This detector-agnostic approach can work with any state-of-the-art detection system without modification, benefiting from advances in object detection while maintaining focus on higher-level event recognition through transparent symbolic reasoning. In summary, our framework bridges the gap between low-level object detection and high-level event understanding through a synergistic combination of symbolic search and LLM reasoning. By discovering interpretable symbolic expressions that operate on detector outputs, we unlock event recognition capabilities without extensive training data or specialized architectures, while maintaining full transparency in the reasoning process.
@@ 4 Experiment
Table 1: Performance of different open set detectors on multiple data sets with or without SymbolicDet module. (AUROC%)
@@
Original +SymbolicDet Original +SymbolicDet Original +SymbolicDet
@@ 4.1 Experimental Setup
To comprehensively evaluate our approach, we conduct experiments across diverse datasets spanning various event detection scenarios, comparing against both traditional methods and architecture variants to demonstrate the efficacy of our LLM-guided symbolic reasoning framework.
@@ 4.1.1 Self-collected Datasets
Multi-Event Dataset is our large-scale collection containing over 110,000 images spanning various event detection scenarios. The dataset comprises over 110,000 images spanning various scenarios including fires and waste incineration (30,954 images), multi-rods fishing (12,000 images), night fishing (97 images), license plate detection (12,487 images), personnel loitering and intrusion (10,788 images), among others. For this study, we specifically focus on the multiple-rod fishing scenario, where we have 15,000 training images with 45,341 detailed bounding box annotations covering persons, fishing rods, tackle bags, and umbrellas. The test set contains 2,283 images, with 1,098 cases showing anomalous multi-rods fishing activities and 1,185 normal cases with single-rod fishing.
@@ 4.1.2 Public Benchmarks
ERA Dataset [42] (Event Recognition in Aerial videos) provides a comprehensive collection of aerial footage covering various event categories. We organize our evaluation around three main event categories: BALL events (327 images) encompassing baseball, soccer, and basketball games; Person_crowded events (352 images) including conflicts, parade protests, and parties; and Sport events (258 images) covering cycling, boating, and racing activities. Additionally, we utilize 347 Non-event images as negative samples, creating a balanced evaluation framework for our method’s discriminative capabilities across different event types.
@@ 4.1.3 Implementation Details
Our implementation integrates three key components: open-vocabulary object detection, LLM-guided symbolic pattern discovery, and Computational Resources.
Open-vocabulary object detector Setup We employ three state-of-the-art multimodal detectors in our experiments: APE [51] (serving as the primary detector), GLIP [27], and YOLO-WORLD [6]. To optimize detection performance, we implement a two-stage prompt generation process. Initially, we leverage LLM to analyze event scenario descriptions and generate comprehensive detection prompts. The LLM generates prompts not only for objects directly associated with event scenarios but also for contextually related non-event objects, ensuring comprehensive coverage of potential scene elements. Based on empirical studies of detector characteristics, we configure different detection thresholds for optimal performance. Considering APE’s characteristically lower threshold nature, we set its minimum detection threshold to 0.05. For GLIP and YOLO-WORLD, we establish a higher threshold of 0.1 to maintain a balance between precision and recall in object detection.
@@
Computational Resources Our experimental framework utilizes a mixed compute infrastructure optimized for different computational demands. Object detection inference is performed on a single NVIDIA RTX 4090 GPU with 24GB memory. Due to our framework’s plug-and-play design, even traditional object detectors can be integrated with minimal resource requirements, making our approach adaptable to various hardware configurations. The symbolic regression component of SymbolicDet runs on Intel(R) Xeon(R) Silver 4214R CPU processors, which are well-suited for the parallel exploration of symbolic search. For LLM reasoning, we utilize qwen-series models. This distributed computational approach ensures efficient processing of our training-free pattern discovery pipeline while maintaining practical performance for real-world applications.
4.2 Main Results
We evaluate our approach from multiple perspectives: effectiveness across different object detection architectures, comparison with fine-tuning approaches, and benchmarking against traditional event detection methods.
Comparison with Different Detection Architectures. We first evaluate our framework using three state-of-the-art open-vocabulary detectors: APE, GLIP, and YOLO-WORLD, which represent diverse architectural choices in both visual and language processing. These detectors employ different visual backbones (VIT [8], Swin-L [37], and YOLOv8) and language models (EVA-CLIP [54], CLIP [45], and BERT [7]). As shown in Table 1, all three detectors achieve strong performance without any fine-tuning, with APE consistently outperforming others across all five anomaly event scenarios. This superior performance of APE can be attributed to its more sophisticated visual-language alignment mechanism and larger pre-training dataset, which enables better transfer of knowledge to anomaly detection tasks. The consistent performance across architecturally diverse models also suggests that our framework’s effectiveness is not tied to specific architectural choices, but rather stems from the fundamental synergy between symbolic reasoning and detection capabilities.
Finding 1: The effectiveness of our training-free framework is architecture-agnostic, with APE’s superior performance likely due to its enhanced visual-language alignment and broader pre-training.
Comparison with Fine-tuning Approaches. To further validate the efficiency of our training-free approach, we conduct comparative experiments with fine-tuned variants on the Helmet-Mac and Multi-rods Fishing datasets. We implement two common fine-tuning strategies: LORA fine-tuning[19] and Prompt tuning [24], representing different levels of parameter adaptation. Results in Table 2 demonstrate that our training-free approach achieves comparable performance to fine-tuned models. This intriguing finding suggests that our training-free approach effectively leverages the model’s general understanding of visual-language relationships. Additionally, the symbolic reasoning component provides a more structured way to capture visual event patterns compared to implicit learning through fine-tuning, achieving similar effectiveness without the computational overhead of parameter adaptation.
Finding 2: Our training-free approach is comparable to fine-tuning methods, possibly due to better preservation of general visual-language understanding and more structured pattern discovery.
Table 2: Performance of different fine-tuned methods and with or without SymbolicDet module. Lora indicates whether to Lora-tuning the APE model. Prompt indicates whether to Prompt-tuning the APE model (AUROC%)
Lora Prompt Our Helmet-Mac Multi-rods Fishing
67.41 66.82
84.11 75.86
67.42 66.82
83.18 (+15.77) 75.16 (+8.34)
95.67 (+11.56) 78.44 (+2.58)
81.62 (+14.2) 76.06 (+9.24)
Comparison with present Methods. To contextualize our approach within the broader landscape of event detection methods, we evaluate on the UCSD Ped2 benchmark and compare against state-of-the-art approaches. As showed in Table 3, our method achieves an impressive 98.7% accuracy without utilizing semantic-level annotations or task-specific fine-tuning, approaching the state-of-the-art performance (99.7%). This minimal performance gap is particularly interesting considering the vast difference in approach complexity. We hypothesize that this effectiveness stems from two factors: first, the pre-trained detectors already possess rich semantic understanding that generalizes well to visual event pattern; second, our symbolic reasoning framework effectively translates this semantic knowledge into explicit detection rules, potentially capturing patterns that are similar to those learned by supervised methods but in a more interpretable manner.
Finding 3: The near-SOTA performance on UCSD Ped2 suggests that combining pre-trained knowledge with symbolic reasoning can effectively match supervised learning capabilities.
Table 3: The overall performance on the UCSD ped2 [56] benchmark.This intuitively reflects that the performance of SymbolicDet (APE) without training is very close to the current SOTA.
@@ 4.2 Main Results
SymbolicDet 98.7
Ablation Studies. We conducted comprehensive ablation studies to evaluate the contribution of each component in our framework. Our analyses examined: (i) the individual impacts of symbolic regression and manual logic, revealing an 18.36% performance improvement with symbolic pattern discovery; (ii) the significant benefits of LLM integration on both accuracy and convergence efficiency; and (iii) the positive correlation between search scale and detection performance across datasets. These experiments not only validate SymbolicDet’s architectural choices but also confirm that its effectiveness stems from the synergistic combination of symbolic reasoning capabilities and LLM-guided semantic understanding. Detailed results, additional visualizations, and in-depth discussion of these ablation studies are provided in the supplementary materials.
5 Conclusion
In this paper, we introduce SymbolicDet, a framework that unlocks event understanding capabilities within standard object detectors through LLM-guided symbolic reasoning. Our approach demonstrates that object detectors contain sufficient visual information for complex event understanding when enhanced with appropriate reasoning mechanisms. Our key contributions include: First, establishing a paradigm that bridges visual perception and symbolic reasoning through evolutionary pattern discovery and LLM guidance, achieving competitive performance with interpretable reasoning. Second, demonstrating an effective training-free framework that eliminates task-specific fine-tuning. Third, contributing two new benchmark datasets for visual event detection research. Our results show that combining pre-trained detectors with explicit symbolic reasoning offers a powerful alternative to specialized, training-intensive approaches while enhancing interpretability and adaptability through human-readable symbolic expressions. Looking forward, while demonstrated in event detection, our approach of enhancing pre-trained visual models with explicit reasoning has broader potential. Future work could extend this framework to relationship detection, behavioral analysis, and intention recognition — further bridging the gap between perception and reasoning in visual understanding systems.
Supplementary Material
6 Ablation Study
@@
Figure 3: Performance on SymbolicDet with or without LLM.
Impact of LLM Integration. Figure 3 illustrates the substantial impact of different LLM integration on both the effectiveness and efficiency of our symbolic pattern discovery process. When examining convergence trajectories across generations, we observe that LLM guidance not only enhances the ultimate detection accuracy but also significantly accelerates the convergence speed of symbolic regression. The analysis compares performance curves with and without LLM guidance, as well as across different LLM scales. Finding: Effective event detection through symbolic reasoning benefits from the complementary strengths of systematic pattern discovery (through evolutionary search) and semantic guidance (through LLM reasoning). The symbolic component provides the expressive framework for capturing complex relationships, while the LLM component contributes domain knowledge and conceptual understanding that steers the search toward meaningful patterns.
Figure 4: Performance on different search scales.
Effect of Search Scale on Performance. To further explore the robustness of our framework, we investigate the effect of varying search scales on event detection accuracy, as depicted in Figure 4. The search scale defines the proportion of samples allocated for constructing the logical search space, with the remainder used for pattern evaluation. Our results reveal a clear pattern: increasing the search scale consistently enhances AUROC performance across both the Helmet and Fishing datasets using APE and GLIP strategies. Notably, in the Helmet dataset, both strategies show a significant improvement, reaching peak performance at the highest search scale of 8.16%. The Fishing dataset demonstrates a similar upward trend, highlighting the benefits of expanding the search space. Finding: The increase in performance with larger search scales underscores the efficacy of our approach in utilizing more extensive logical reasoning. The findings suggest that even without traditional fine-tuning, enlarging the search space enables the framework to uncover more accurate and interpretable patterns. This scalability evidences the flexibility and potency of SymbolicDet in capitalizing on the latent potential of standard object detectors, reinforcing its applicability across diverse scenarios.
Figure 5: Illustration of the application of symbolic pattern detection in sports environments, showcasing how logical expressions can be used to identify and categorize complex sports scenarios.